# Feynman Graph Integrals on Kähler Manifolds

### v2

@@ 3.1 Regular expressions and its filtration
Proposition 3.8 shows that the filtration $F_{*}A(N(U,r);p)$ is independent of the choice of coordinates. This allows us to define the filtration globally on complex manifold. The idea of introducing scaling actions and filtrations is motivated by Getzler’s rescaling, which plays a crucial role in the study of heat kernel expansions and index theorems. See [4, 15] for further details.
@@ 4 Regularities of Propagator in Schwinger Space
The linear space of regular expressions $S_{\mathrm{reg}}\subset\Omega^{*}\Big{(}(0,\infty);\Gamma\big{(}{\widetilde{E}}\boxtimes{\widetilde{E}}\big{)}\Big{)}$ is generated by differential forms
$w\in\Omega^{*}\Big{(}(0,\infty);\Gamma\big{(}{\widetilde{E}}\boxtimes{\widetilde{E}}\big{)}\Big{)}$,
@@ 4.1 Heat kernel expansion
The second statement follows directly from [30, Theorem 1.1]. Here we briefly explain the proof of the first statement.
@@
The heat kernel coefficients $u_{i}$ can be computed recursively by solving ODEs, as described in [4, Chapter 2]. Since both $(M,g)$ and the bundle $E\to M$ are real analytic, and the initial condition is given by (4.5), it follows from standard theory of ODEs that each $u_{i}$ is analytic.
@@ 4.2 Getzler’s rescaling
Then one can see that $C_{l},l\leq 0$ satisfies the following equation:
@@
Let $q_{t}:=e^{-\frac{|\tilde{{\mathbf{z}}}|^{2}}{2t}}$. Restricted on $U\times\{p\}$,
$L\big{(}q_{t}C_{l}\big{)}=0,\,l\leq 0.$
Here, $L$ denotes the operator defined in (4.10).
@@
Since, by the construction above, $\delta_{\epsilon}C_{l}=C_{l}\epsilon^{l}$, and $\delta_{\epsilon}C_{1}=O(\epsilon)$, we have
$\delta_{\epsilon}\Big{(}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}=\sum_{l=-n}^{0}C_{l}\epsilon^{l}+O(\epsilon).$ (4.11)
Following the standard theory of heat kernel expansions (see, e.g., [4, Chapter 2]), we have
$\left(\partial_{t}+\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)\Big{(}e^{-\frac{\rho^{2}}{2t}}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}=e^{-\frac{\rho^{2}}{2t}}\left(\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)u_{n}.$ (4.12)
It follows from (A.1) that, restricted to $U\times\{p\}$,
@@
Restricted to $U\times\{p\}$, and using 4.7, together with (4.11)–(4.13), we compute:
$L\bigg{(}q_{t}\Big{(}\sum_{l=-n}^{0}C_{l}\epsilon^{l}\Big{)}\bigg{)}$ $=L\Big{(}\delta_{\epsilon}e^{-\frac{\rho^{2}}{2t}}+O(\epsilon)\Big{)}\bigg{(}\delta_{\epsilon}\Big{(}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}+O(\epsilon)\bigg{)}$
$=L\Big{(}\delta_{\epsilon}e^{-\frac{\rho^{2}}{2t}}\Big{)}\delta_{\epsilon}\Big{(}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}+O(\epsilon)$
$=\left(\epsilon\,\delta_{\epsilon}\left(\partial_{t}+\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)\delta_{\epsilon}^{-1}\right)\delta_{\epsilon}\Big{(}e^{-\frac{\rho^{2}}{2t}}\sum_{i=0}^{n}t^{i-n}u_{i}\Big{)}+o(1)$
$=\epsilon\,\delta_{\epsilon}\Big{(}e^{-\frac{\rho^{2}}{2t}}\left(\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)u_{n}\Big{)}+o(1)$
$=o(1).$
Examining the order in $\epsilon$, the proposition follows.
@@
As a result, we have:
Corollary 4.10.
Restricted on $U\times\{p\}$, we have $C_{l}=0$ if $l<0$.
Proof.
Let $q_{t}:=e^{-\frac{|\tilde{{\mathbf{z}}}|^{2}}{2t}}$. By 4.8,
$L\left(q_{t}C_{l}\right)=0,l\leq 0$
Note that $C_{l}=\sum_{i=0}^{n}t^{i-n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=l+n-i}\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{k}}d{\bar{\tilde{{\mathbf{z}}}}}^{\mathbf{j}}\wedge d{\bar{\mathbf{w}}}^{\mathbf{l}}$, the coefficient of $t^{-n}$ is
However, by (4.2), associated with the frame $(U,{\mathbf{z}},{\mathbf{e}})$,
$u_{0}|_{\{p\}\times\{p\}}$ $=\omega(z,w)d^{n}(\bar{\mathbf{z}}-\bar{\mathbf{w}})|_{\{0\}\times\{0\}}.$ (4.14)
Thus $C_{l}^{0}|_{\{p\}\times\{p\}}=0$ for $l<0$ (since in this case, the multi-index $\mathbf{j}$} in the summation defining $C_{l}^{0}$ satisfies $|{\mathbf{j}}|\leq n+l<n$). Therefore, by 4.9, it follows that $C_{l}=0$ for all $l<0$. ∎
By 4.10 and 4.6, restricted on $U\times\{p\}$,
@@
It follows from 4.11 that
@@
By 4.11 and note that $\iota_{\frac{\partial}{\partial{\bar{z}}^{i}}}{d\bar{\mathbf{w}}}^{\mathbf{l}}=0$, we compute
@@
Next, proceeding as in 4.12, we have
@@
By (2) in 4.11 and the fact that $\frac{b_{0}}{t}\in{S}_{{\mathrm{reg}}}$ (see 4.5), we can set:
@@
Together with 4.12, the result follows. ∎
@@
${\mathcal{M}}=\sum_{i,j}\Big{(}\frac{1}{2t}\frac{\partial\rho^{2}}{\partial z_{i}}(g^{-1})^{i\bar{j}}\iota_{\frac{\partial}{\partial{\bar{z}}_{j}}}-(g^{-1})^{i\bar{j}}\iota_{\frac{\partial}{\partial{\bar{z}}_{j}}}\nabla^{{\widetilde{E}}}_{\frac{\partial}{\partial z_{i}}}\Big{)}$. (4.22)
@@
where the first equality follows from the fact that $(dt)^{2}=0$, along with (4.19) and 4.14; the second equality follows from 4.12.
Note that $P_{t}=-dt\wedge(\bar{\partial}_{{\widetilde{E}}}^{*}\otimes\mathrm{id})H_{t}+H_{t}$, by 4.13 and the computation above, (4.23) follows. ∎
@@
The identity (4.24) follows directly from 4.15 and 4.16. Therefore, 4.3 is proved.
@@
The proof of 4.7 follows from the standard argument of Getzler’s rescaling [15, 4]; we include it here for completeness.
@@
Note that on $U\times\{p\}$, where ${\tilde{z}}=z$, ${\bar{\tilde{z}}}={\bar{z}}$, we have:
@@ Appendix A Existence of Canonical Local Coordinates
The proof can be found in [22, 32]. ∎
@@
We recall the Eikonal equation on Kähler manifolds (see [29, Example 3.2]):
@@ Appendix B Cauchy Principal Value
In this appendix, we will introduce the Cauchy principal value for the reader’s convenience. We will mainly follow [18], with extra attention to logarithmic terms.
@@
Given a differential form $\alpha\in\Omega^{*,*}\left(\left(\mathbb{C}^{*}\right)^{m}\times\mathbb{C}^{n}\right)$. We say $\alpha$ is a differential form on $\mathbb{C}^{m+n}$ with divisorial type singularities of $({\mathbf{i}},{\mathbf{k}})$-type along the principal divisor: $f=z_{1}z_{2}\cdots z_{m}$, if there exist $\beta\in\Omega^{*,*}\left(\mathbb{C}^{m+n}\right)$, ${\mathbf{i}}\in\mathbb{Z}_{\geq 0}^{m}$ and $\bf\in\mathbb{Z}_{\geq 0}^{m}$, such that
@@
Let $B:=\{{\mathbf{z}}\in\mathbb{C}^{m+n}:\sup_{1\leq i\leq m+n}|z_{i}|\leq 1\}.$
@@
Similar to [18, Lemma 6.4], we have
@@
Let $\phi\in L^{\infty}(\mathbb{C}^{m+n})$, then for any ${\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m},{\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$, any $i\in\{1,2,\cdots,m\}$,
@@
Let $\mathbf{k}=(k_{1},\dots,k_{m})$ and $\mathbf{j}=(j_{1},\dots,j_{m})$, and denote $|\mathbf{j}|:=j_{1}+\dots+j_{m}$.
@@
$(\theta,{\mathbf{z}}_{\hat{1}})\mapsto\Bigg{(}\bigg{(}\frac{\delta}{\prod_{s=2}^{m}\rho_{s}^{j_{s}}}\bigg{)}^{1/j_{1}}e^{i\theta},{\mathbf{z}}_{\hat{1}}\Bigg{)}$,
@@
Similar to [18, Lemma 6.3], we have
@@
    $\int_{B_{\delta}^{\mathbf{j}}}z_{l}^{r}\bar{z}_{l}^{s}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=0;$
@@
    $\int_{S_{\delta}^{\mathbf{j}}}z_{l}^{r}\bar{z}_{l}^{s}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0$.
@@
For each ${\mathbf{z}}_{\hat{l}}\in B_{\delta,\hat{l}}^{\mathbf{j}}$, define
$B({\mathbf{z}}_{\hat{l}},\delta)=\left\{z_{l}\in\mathbb{C}:\delta/|{\mathbf{z}}_{\hat{l}}^{{{\mathbf{j}}}_{\hat{l}}}|<|z_{l}^{j_{l}}|<1\right\},$
The partial integrals
$I({\mathbf{z}}_{\hat{l}},\delta)$ $=\int_{B({\mathbf{z}}_{\hat{l}},\delta)}z_{l}^{r-i_{l}}\bar{z}_{l}^{s}\big{(}\ln(|z_{l}|^{2})\big{)}^{k_{l}}dz_{l}\wedge d\bar{z}_{l}$
$=\int_{\left(\delta/|{\mathbf{z}}_{\hat{l}}^{{\mathbf{j}}_{\hat{l}}}|\right)^{\frac{1}{j_{l}}}}^{1}\rho^{r-i_{l}+s+1}\big{(}\ln(\rho^{2})\big{)}^{k_{l}}d\rho\int_{0}^{2\pi}e^{i\theta\left(r-i_{l}-s\right)}d\theta$
are zero for all ${\mathbf{z}}_{\hat{l}}\in B_{\delta,\hat{l}}^{\mathbf{j}}$, because $r-s\leq r+s<i_{l}$. Then the integral in (1) is zero, since it is equal to
$\int_{B_{\delta,\hat{l}}^{{\mathbf{j}}}}I({\mathbf{z}}_{\hat{l}},\delta)\cdot{\mathbf{z}}_{\hat{l}}^{-{\mathbf{i}}_{\hat{l}}}\ln({\mathbf{z}}_{\hat{l}},{\mathbf{k}}_{\hat{l}})gd{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}$.
@@
which are all zero if $s\geq 0$, since in such case $r-i_{l}-s-1\leq r-i_{l}+s-1<0$. We deduce that the integral in (2) is zero. ∎
Similar to [18, Lemma 6.2],
@@
$k({\mathbf{z}},\bar{{\mathbf{z}}})=\sum_{r+s<i_{l}}z_{l}^{r}\bar{z}_{l}^{s}g_{r,s}^{l}({\mathbf{z}}_{\hat{l}},\bar{{\mathbf{z}}}_{\hat{l}})+K^{l}({\mathbf{z}},\bar{{\mathbf{z}}})$
@@
The lemma is nearly identical to [18, Lemma 6.2], except that it does not involve summation over $l=1,\dots,m$. Consequently, the expression $g_{r,s}^{l}$ differs slightly from that in [18, Lemma 6.2].
@@
Proceeding as in [18, Proposition 6.5], using B.4-B.6, we have:
@@
For any ${\mathbf{i}},{\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m},{\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$ and $g\in C^{\infty}\left(\mathbb{C}^{m+n}\right)$, the limit
$\lim_{\delta\rightarrow 0}\int_{B_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$ (B.5)
exists. Moreover,
$\lim_{\delta\rightarrow 0}\int_{S_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0.$ (B.6)
@@
Let $f$ be a never vanishing smooth function on $\bar{B}\subset\mathbb{C}^{m+n}$, and let ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m},{\mathbf{i}}\in\mathbb{Z}_{\geq 0}^{m}$. Denote
$S_{\delta}^{\mathbf{j}}(f)=\big{\{}{\mathbf{z}}\in B:|f{\mathbf{z}}^{\mathbf{j}}|=\delta\big{\}}$
$B_{\delta}^{\mathbf{j}}(f)=\big{\{}{\mathbf{z}}\in B:|f{\mathbf{z}}^{\mathbf{j}}|>\delta\big{\}}$
Consider $B_{\delta}^{\mathbf{j}}(f)$ with its canonical complex orientation and $S_{\delta}^{\mathbf{j}}(f)$ with the orientation such that $\partial B_{\delta}^{\mathbf{j}}(f)=-S_{\delta}^{\mathbf{j}}(f)$.
There exists an open neighborhood $\tilde{B}\subset B$ of the origin $0\in\mathbb{C}^{m+n}$ such that, if $\tilde{S}_{\delta}^{\mathbf{j}}:=S_{\delta}^{\mathbf{j}}(f)\cap\tilde{B}$ and $\tilde{B}_{\delta}^{\mathbf{j}}:=B_{\delta}^{\mathbf{j}}(f)\cap\tilde{B}$, then
$\lim_{\delta\rightarrow 0}\int_{\tilde{S}_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}=0$ (B.7)
for any $g\in C^{\infty}(\bar{B})$, and
$\lim_{\delta\rightarrow 0}\int_{\tilde{B}_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\lim_{\delta\rightarrow 0}\int_{B_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$ (B.8)
[…]

### v3

@@ 3.1 Regular expressions and its filtration
Proposition 3.8 shows that the filtration $F_{*}A(N(U,r);p)$ is independent of the choice of coordinates. This allows us to define the filtration globally on complex manifold. The idea of introducing scaling actions and filtrations is motivated by Getzler’s rescaling, which plays a crucial role in the study of heat kernel expansions and index theorems. See [berline2003heat, getzler1986short] for further details.
@@ 4 Regularities of Propagator in Schwinger Space
The linear space of regular expressions $)S_{\mathrm{reg}}\subset\Omega^{*}\Big((0,\infty);\Gamma\big({\widetilde{E}}\boxtimes{\widetilde{E}}\big$$\Big)$ is generated by differential forms
$),w\in\Omega^{*}\Big((0,\infty);\Gamma\big({\widetilde{E}}\boxtimes{\widetilde{E}}\big$$\Big)$,
@@ 4.1 Heat kernel expansion
The second statement follows directly from [ludewig2019strong, Theorem 1.1]. Here we briefly explain the proof of the first statement.
@@
The heat kernel coefficients $u_{i}$ can be computed recursively by solving ODEs, as described in [berline2003heat, Chapter 2]. Since both $(M,g)$ and the bundle $E\to M$ are real analytic, and the initial condition is given by (4.5), it follows from standard theory of ODEs that each $u_{i}$ is analytic.
@@ 4.2 Getzler’s rescaling
More explicitly,
$L=\epsilon\delta_{\epsilon}\left(\partial_{t}+\Delta_{\widetilde{E}}\otimes\mathrm{id}\right)\delta_{\epsilon}^{-1}+O(\epsilon)$
We will show that
@@
Let $q_{t}:=e^{-|\tilde{{\mathbf{z}}}|^{2}/2t}$. Then, restricted to $U\times\{p\}$,
  1. (a)
    Let $L$ denote the operator defined in (4.10). Then
    $L\big(q_{t}C_{l}\big)=0,\quad l\leq 0.$
  2. (b)
    We have
    $C_{l}=0\quad\text{for }l<0.$
@@
We first treat the case l=nl=-n. By the construction above, $\delta_{\epsilon}C_{l}=C_{l}\epsilon^{l}$ and $\delta_{\epsilon}C_{1}=O(\epsilon)$. Hence
$\delta_{\epsilon}\!\left(\sum_{i=0}^{n}t^{i-n}u_{i}\right)=\sum_{l=-n}^{0}C_{l}\epsilon^{l}+O(\epsilon).$ (4.11)
From the standard heat kernel expansion (see, e.g., [berline2003heat, Chap. 2]),
$\bigl(\partial_{t}+\Delta_{\widetilde{E}}\otimes\mathrm{id}\bigr)\!\left(e^{-\rho^{2}/2t}\!\sum_{i=0}^{n}t^{i-n}u_{i}\right)=e^{-\rho^{2}/2t}\bigl(\Delta_{\widetilde{E}}\otimes\mathrm{id}\bigr)u_{n}.$ (4.12)
By (A.1), on $U\times\{p\}$,
@@
Restricting to $U\times\{p\}$ and using 4.7 together with (4.11)–(4.13),
$\ \ \ \ L\!\left(q_{t}\!\sum_{l=-n}^{0}C_{l}\epsilon^{l}\right)=L\!\left(\delta_{\epsilon}e^{-\rho^{2}/2t}+O(\epsilon)\right)\left(\delta_{\epsilon}\!\sum_{i=0}^{n}t^{i-n}u_{i}+O(\epsilon)\right)$ (4.14)
$=L(\delta_{\epsilon}e^{-\rho^{2}/2t})\,\delta_{\epsilon}\!\sum_{i=0}^{n}t^{i-n}u_{i}+O(\epsilon^{-n+1})$
$=\epsilon\,\delta_{\epsilon}\bigl((\partial_{t}+\Delta_{\widetilde{E}}\otimes\mathrm{id})\delta_{\epsilon}^{-1}\bigr)\delta_{\epsilon}\!\left(e^{-\rho^{2}/2t}\!\sum_{i=0}^{n}t^{i-n}u_{i}\right)+O(\epsilon^{-n+1})$
$=\epsilon\,\delta_{\epsilon}\!\left(e^{-\rho^{2}/2t}(\Delta_{\widetilde{E}}\otimes\mathrm{id})u_{n}\right)+O(\epsilon^{-n+1})=O(\epsilon^{-n+1}).$
Comparing powers of $\epsilon$ gives
$L(q_{t}C_{-n})=0.$
To show (b) for l=nl=-n, write
$C_{l}=\sum_{i=0}^{n}t^{i-n}\sum_{|{\mathbf{j}}|+|{\mathbf{k}}|=l+n-i}\;\sum_{{\mathbf{l}}}b^{i}_{{\mathbf{k}},{\mathbf{j}},{\mathbf{l}}}\,{\bar{\tilde{{\mathbf{z}}}}}^{{\mathbf{k}}}d{\bar{\tilde{{\mathbf{z}}}}}^{{\mathbf{j}}}\wedge d\bar{{\mathbf{w}}}^{{\mathbf{l}}}.$
The coefficient of $t^{-n}$ is
By (4.2), in the frame $(U,{\mathbf{z}},{\mathbf{e}})$,
$u_{0}|_{\{p\}\times\{p\}}=\omega({\mathbf{z}},{\mathbf{w}})\,d^{n}(\bar{{\mathbf{z}}}-\bar{{\mathbf{w}}})|_{\{0\}\times\{0\}}.$
Thus $C_{l}^{0}|_{\{p\}\times\{p\}}=0$ for $l<0$, since then $|{\mathbf{j}}|\leq n+l<n$. By 4.9 below, this implies $C_{-n}=0$.
With $C_{-n}=0$, (4.14) improves to
$L\!\left(q_{t}\sum_{l=-n+1}^{0}C_{l}\epsilon^{l}\right)=O(\epsilon^{-n+2}).$
Repeating the argument yields $L(q_{t}C_{-n+1})=0$ and therefore $C_{-n+1}=0$. Proceeding inductively gives $C_{l}=0$ for all $l<0$ and $L(q_{t}C_{l})=0$ for all $l\leq 0$.
@@
By 4.8 and 4.6, restricted on $U\times\{p\}$,
@@
It follows from 4.10 that
@@
By 4.10 and note that $\iota_{\frac{\partial}{\partial{\bar{z}}^{i}}}{d\bar{\mathbf{w}}}^{\mathbf{l}}=0$, we compute
@@
Next, proceeding as in 4.11, we have
@@
By (2) in 4.10 and the fact that $\frac{b_{0}}{t}\in{S}_{{\mathrm{reg}}}$ (see 4.5), we can set:
@@
Together with 4.11, the result follows. ∎
@@
[…]
@@
where the first equality follows from the fact that $(dt)^{2}=0$, along with (4.19) and 4.13; the second equality follows from 4.11.
Note that $P_{t}=-dt\wedge(\bar{\partial}_{{\widetilde{E}}}^{*}\otimes\mathrm{id})H_{t}+H_{t}$, by 4.12 and the computation above, (4.23) follows. ∎
@@
The identity (4.24) follows directly from 4.14 and 4.15. Therefore, 4.3 is proved.
@@
The proof of 4.7 follows from the standard argument of Getzler’s rescaling [getzler1986short, berline2003heat]; we include it here for completeness.
@@
Note that on $U\times\{p\}$, where $\tilde{{\mathbf{z}}}={\mathbf{z}}$, $\bar{\tilde{{\mathbf{z}}}}=\bar{{\mathbf{z}}}$, we have:
@@ Appendix A Existence of Canonical Local Coordinates
The proof can be found in [Kontsevich1994-1995, ruan1996canonical]. ∎
@@
We recall the Eikonal equation on Kähler manifolds (see [doi:10.1142/S0129055X24500478, Example 3.2]):
@@ Appendix B Cauchy Principal Value
In this appendix, we will introduce the Cauchy principal value for the reader’s convenience. We will mainly follow [herrera1971residues], with extra attention to logarithmic terms.
@@
Given a differential form $\alpha\in\Omega^{*,*}\left(\left(\mathbb{C}^{*}\right)^{m}\times\mathbb{C}^{n}\right)$. We say $\alpha$ is a differential form on $\mathbb{C}^{m+n}$ with divisorial type singularities of $({\mathbf{i}},{\mathbf{k}})$-type along the principal divisor: $f=z_{1}z_{2}\cdots z_{m}$, if there exist $\beta\in\Omega^{*,*}\left(\mathbb{C}^{m+n}\right)$, ${\mathbf{i}}\in\mathbb{Z}_{\geq 0}^{m}$ and ${\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m}$, such that
@@
Let
$B:=\{{\mathbf{z}}\in\mathbb{C}^{m+n}:\sup_{1\leq i\leq m+n}|z_{i}|<1\}.$ (B.2)
Let $\mathbf{k}\in\mathbb{Z}_{\geq 0}^{m}$, with components $(k_{1},\dots,k_{m})$, and denote
$|\mathbf{k}|:=k_{1}+\dots+k_{m}.$
@@
Similar to [herrera1971residues, Lemma 6.4], we have
@@
Let $\phi\in C\big((\mathbb{C}^{*})^{m}\times\mathbb{C}^{n}\big)$ be bounded. Then for any ${\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m}$, ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$, and any $i\in\{1,2,\dots,m\}$,
@@
[…]
@@
Similar to [herrera1971residues, Lemma 6.3], we have
@@
    $\int_{S_{\delta}^{\mathbf{j}}}z_{l}^{r}\bar{z}_{l}^{s}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0$;
@@
    If $s\geq 1$, then $\int_{S_{\delta}^{\mathbf{j}}}z_{l}^{r}\bar{z}_{l}^{s}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0$.
@@
which are all zero if $s\geq 0$, since in such case $r-i_{l}-s-1\leq r-i_{l}+s-1<0$. We deduce that the integral in (1) is zero.
For (2), just note that if $s\geq 1$, $r-i_{l}-s+1\leq r-i_{1}+s-2+1<0$. ∎
Similar to [herrera1971residues, Lemma 6.2],
@@
$k({\mathbf{z}},\bar{{\mathbf{z}}})=\sum_{r+s<i_{l}}z_{l}^{r}\bar{z}_{l}^{s}g_{r,s}^{l}({\mathbf{z}}_{\hat{l}},\bar{{\mathbf{z}}}_{\hat{l}})+K^{l}({\mathbf{z}},\bar{{\mathbf{z}}}),$
@@
The lemma is nearly identical to [herrera1971residues, Lemma 6.2], except that it does not involve summation over $l=1,\dots,m$. Consequently, the expression $g_{r,s}^{l}$ differs slightly from that in [herrera1971residues, Lemma 6.2].
@@
Proceeding as in [herrera1971residues, Proposition 6.5], using B.4-B.6, we have:
@@
For any ${\mathbf{i}},{\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m},{\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$ and $g\in C^{\infty}\left(\mathbb{C}^{m+n}\right)$, we have
$\lim_{\delta\rightarrow 0}\int_{S_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})gd\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0$ (B.6)
and
$\lim_{\delta\rightarrow 0}\int_{S_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\bar{z}_{l}gd{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0.$ (B.7)
Thus, we obtain the following.
Corollary B.8.
For any ${\mathbf{i}},{\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m}$, ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$, and $g\in C_{c}^{\infty}(\mathbb{C}^{m+n})$, we have
$\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\,\ln({\mathbf{z}},{\mathbf{k}})\,g\;d\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0,$ (B.8)
and
$\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\,\ln({\mathbf{z}},{\mathbf{k}})\,\bar{z}_{l}g\;dz_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=0.$ (B.9)
Proof.
We may assume $g\in C_{c}^{\infty}(B)$. Indeed, if not, there exists $N>1$ such that $\mathrm{supp}(g)\subset B_{N}:=\{{\mathbf{z}}\in\mathbb{C}^{m+n}:\sup_{1\leq j\leq m+n}|z_{j}|<N\}$. In this case, one verifies that B.7 continues to hold when the set $B$ is replaced by $B_{N}$.
Now assume $g\in C_{c}^{\infty}(B)$. Then
$\int_{|{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\,\ln({\mathbf{z}},{\mathbf{k}})\,g\;d\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}=\int_{S_{\delta}}{\mathbf{z}}^{-{\mathbf{i}}}\,\ln({\mathbf{z}},{\mathbf{k}})\,g\;d\bar{z}_{l}\wedge d{\mathbf{z}}_{\hat{l}}\wedge d\bar{{\mathbf{z}}}_{\hat{l}}$,
so (B.8) follows directly. The argument for (B.9) is analogous. ∎
@@
Let $f$ be a real nowhere vanishing smooth function on $\mathbb{C}^{m+n}$, and let ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$, ${\mathbf{i}},{\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m}$. Then, for any $g\in C_{c}^{\infty}(\mathbb{C}^{m+n})$, we have
$\lim_{\delta\to 0}\int_{|f{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\,\ln({\mathbf{z}},{\mathbf{k}})\,g\;d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}=0.$ (B.10)